OpenStudy (anonymous):

what is the inntegral x^2 sec^2 x^3 dx? what is the integral of x csc^2 x ^2dx?

OpenStudy (anonymous):

\[\int x^2\sec^2\left(x^3\right)\,dx\]\[\int x\csc^2\left(x^2\right)\,dx\]To clarify. Also, try \(u=x^3\) and \(u=x^2\) respectively.